# Jinhang Chai     Jianqing Fan

### v1

@@ B.1 Proof sketch
    Let $g:=\sqrt{\frac{p}{\mu r}}|\lambda|$, by consider separately the cases of $r=1$ and $r\geq 2$, it is straightforward to show
    $(r-1)\cdot\big{(}\frac{1}{2r}\sqrt{\frac{p}{\mu r}}|\lambda|\big{)}^{2}\leq\frac{g^{2}}{16}$
    Putting together we get
    $\|V^{\top}Re_{1}\|^{2}$ $\leq(1-\frac{g}{4})^{2}+\frac{g^{2}}{16}$
    $\leq 1-\frac{g}{2}+\frac{g^{2}}{8}\leq 1-\frac{g}{4}$ (49)
@@
    Hence by the same technique as (37), and by (B.4), we have
@@
    $\geq 1-\sqrt{1-\frac{g}{4}}$
    Further by the definition of $g$ and the $\mu$-incoherence condition,
    $1-\sqrt{1-\frac{g}{4}}$ $\geq\frac{g}{8}=\frac{1}{8}\sqrt{\frac{p}{\mu r}}|\lambda|\geq\frac{a^{2}|\lambda|^{2}}{32r^{2}}\frac{p}{\mu r}$
    Therefore, it holds that $r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{a^{2}|\lambda|^{2}}{32r^{2}}\frac{p}{\mu r}$.
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq c\frac{a^{2}|\lambda|^{2}}{r^{2}}\frac{p}{\mu r}=c\frac{p}{\mu r^{3}}a^{2}(r_{1}-v_{1})^{2}$
B.5 Proof of Lemma B.5
@@
We argue that there is no loss of generality in assuming $u=ae_{1}$ and $l=be_{1}$. To see that, notice that the lemma is invariant up to joint rotation. Speicifically, let
$\widetilde{U}=US,\widetilde{V}=VS^{\top},\widetilde{L}=LT,\widetilde{R}=RT^{\top}$
@@
where $S,T\in\mathcal{O}_{r}$ are arbitrary orthogonal matrices, then
$\langle U^{\top}L,V^{\top}R\rangle=\langle\widetilde{U}^{\top}\widetilde{L},\widetilde{V}^{\top}\widetilde{R}\rangle,u^{\top}v=\widetilde{u}^{\top}\widetilde{v},l^{\top}r=\widetilde{l}^{\top}\widetilde{r}$
In addition, the incoherence parameter $\mu$ stays unchanged after the procedure. Thus we can rotate these matrices such that the first row of $U$ and $L$ align with $e_{1}$ while preserving the lemma statement, i.e., $u=ae_{1}$ and $l=be_{1}$. Furthermore, we can assume $b\geq a\geq 0$, as $\langle U^{\top}L,V^{\top}R\rangle=\langle L^{\top}U,R^{\top}V\rangle$ and we can swap the order of $U$ and $L$, $V$ and $R$.
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq c\frac{p}{\mu r^{3}}\cdot(av_{1}-br_{1})^{2}$
@@
Now by Lemma B.3 we have
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{8\mu r}\cdot r_{1}^{2}(a-b)^{2}\geq\frac{p}{32\mu r^{3}}r_{1}^{2}(a-b)^{2}.$
By Lemma B.4 we have
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{32\mu r^{3}}a^{2}(r_{1}-v_{1})^{2}.$
@@
$r-\langle U^{\top}L,V^{\top}R\rangle$ $\geq\frac{1}{2}\frac{p}{64\mu r^{3}}\big{(}r_{1}^{2}(a-b)^{2}+a^{2}(r_{1}-v_{1})^{2}\big{)}$
$\geq\frac{p}{256\mu r^{3}}(av_{1}-br_{1})^{2}$
@@
In fact, using Lemma B.2 with $r=1$. We get
$U^{\top}L\geq 1-\frac{(u_{1}-l_{1})^{2}}{2}$
@@
$\geq 1-\sqrt{1-\frac{(u_{1}-l_{1})^{2}}{2}}\geq\frac{(u_{1}-l_{1})^{2}}{4}$
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{(u_{1}-l_{1})^{2}}{4}\geq\frac{p}{\mu r}\frac{v_{1}^{2}(u_{1}-l_{1})^{2}}{4}$
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{\mu r}\frac{l_{1}^{2}(v_{1}-r_{1})^{2}}{4}$
@@
$\geq\frac{p}{\mu r^{3}}\frac{(u_{1}v_{1}-l_{1}r_{1})^{2}}{16}$
B.6 Proof of Lemma B.1
We postpone the technical Lemma B.6 to the next subsection and state the proof of the key separation lemma.
Proof of lemma B.1.
@@
$=\sum_{i=1}^{r}\sigma_{i}^{2}+\sum_{i=1}^{r}\lambda_{i}^{2}-2\operatorname{Tr}(R^{\top}V\Sigma U^{\top}L\Lambda)$
@@
By using the Lemma B.5, it’s not hard to see that $X\in\Omega$ where $\Omega$ is defined in (31) with
$\gamma_{k}=c\frac{p}{\mu r^{3}}\big{[}\sum_{i=1}^{k}(u_{i}v_{i}-l_{i}r_{i})\big{]}^{2}.$ (50)
@@
$\gamma_{k}\leq c\frac{p}{\mu r^{3}}2r\frac{\mu r}{p}=\frac{2c}{r}<\frac{1}{3}$
We can apply Lemma B.6 and so
@@
$\geq\sum_{i=1}^{r}\frac{1}{3}(\sigma_{i}-\lambda_{i})^{2}+\sum_{i=1}^{r-1}\gamma_{i}\{(\sigma_{i}-\lambda_{i+1})^{2}+(\lambda_{i}-\sigma_{i+1})^{2}\}+\gamma_{r}(\sigma_{r}^{2}+\lambda_{r}^{2})$ (51)
where $A$ and $X_{0}$ are defined in Lemma B.6. the second equality follows from direct calculation and we let $\gamma_{0}$ for convenience.
On the other hand, let $\phi_{i}=\left\{\begin{aligned} &\sigma_{i},\quad\text{if}\ 2|i\\ &\lambda_{i},\quad\text{if}\ 2\nmid i\end{aligned}\right.$, we have by Cauchy’s inequality,
$\Delta_{1,1}^{2}=(P_{1,1}-Q_{1,1})^{2}$ $=\Big{[}\sum_{i=1}^{r}\big{(}\sigma_{i}u_{i}v_{i}-\lambda_{i}l_{i}r_{i}\big{)}\Big{]}^{2}$ (52)
$\leq 2\Big{[}\sum_{i=1}^{r}\big{(}\phi_{i}u_{i}v_{i}-\phi_{i}l_{i}r_{i}\big{)}\Big{]}^{2}+2r\sum_{i=1}^{r}(\sigma_{i}-\lambda_{i})^{2}(\frac{\mu r}{p})^{2}$ (53)
@@
$\Big{[}\sum_{i=1}^{r}\big{(}\phi_{i}u_{i}v_{i}-\phi_{i}l_{i}r_{i}\big{)}\Big{]}^{2}$ $=r\sum_{k=1}^{r}(\phi_{k}-\phi_{k+1})^{2}\Big{[}\sum_{i=1}^{k}(u_{i}v_{i}-l_{i}r_{i})\Big{]}^{2}$
$=r\frac{\mu r^{3}}{cp}\sum_{k=1}^{r}(\phi_{k}-\phi_{k+1})^{2}\gamma_{k}$
$\leq\frac{\mu r^{4}}{cp}\sum_{i=1}^{r-1}\{(\sigma_{i}-\lambda_{i+1})^{2}+(\lambda_{i}-\sigma_{i+1})^{2}\}^{2}\gamma_{i}+\frac{\mu r^{4}}{cp}(\sigma_{r}^{2}+\lambda_{r}^{2})\gamma_{r}$ (54)
where the first inequality uses Cauchy’s inequality, the first equality uses (50) and the last inequality follows from the definition of $\phi$.
Therefore, choosing $c$ sufficiently large and combining (51),(53),(54) gives
$\Delta_{1,1}^{2}$ $\leq\frac{2\mu r^{4}}{cp}\sum_{i=1}^{r-1}\{(\sigma_{i}-\lambda_{i+1})^{2}+(\lambda_{i}-\sigma_{i+1})^{2}\}^{2}\gamma_{i}+\frac{2\mu r^{4}}{cp}(\sigma_{r}^{2}+\lambda_{r}^{2})\gamma_{r}+2r\sum_{i=1}^{r}(\sigma_{i}-\lambda_{i})^{2}(\frac{\mu r}{p})^{2}$
$\leq\frac{2\mu r^{4}}{cp}\left(\sum_{i=1}^{r}\frac{1}{3}(\sigma_{i}-\lambda_{i})^{2}+\sum_{i=1}^{r-1}\gamma_{i}\{(\sigma_{i}-\lambda_{i+1})^{2}+(\lambda_{i}-\sigma_{i+1})^{2}\}+\gamma_{r}(\sigma_{r}^{2}+\lambda_{r}^{2})\right)$
$\leq\frac{2\mu r^{4}}{cp}\|\Delta\|_{F}^{2}$
where the second inequality utilizes implicitly $pr\geq 3c$ which is obvious when $c\leq\frac{1}{3}$. Set $\tilde{c}=\frac{2}{c}$ finishes the proof.
B.7 Proof of Lemma B.6
Proof of lemma B.6.
@@
    It is straightforward to observe that $\Omega$ is a polyhedron and the problem is LP.
@@
    Because the $\epsilon$-ball $\{X\big{|}\|X\|_{F}\leq\epsilon\}\subset\Omega$ if we choose $\epsilon$ small enough, we can draw that $\Omega$ is non-degenerate.
@@
Finally, we prove that every feasible direction from $X_{0}$ only decreases the objective, hence finishing the proof.
To obtain that, let $\Delta^{0}$ be an arbitrary feasible direction, and we want to prove $\langle A,\Delta^{0}\rangle\leq 0$ where $A$ is defined in (30). In later proof, we construct an operation and a property such that the initial $\Delta^{0}$ satisfies the property, the operation increases the objective function $\langle A,\Delta^{0}\rangle$ while preserving the prescribed property. Finally, the $X_{0}$ will end up being a zero matrix and $\langle A,\Delta^{0}\rangle=0$.
The property $\mathcal{P}$ defined for a $X\in\mathbb{R}^{r\times r}$ is sum of every upper left block of $X$ is greater or equal to 0, that is,
$\sum_{i=1}^{k_{1}}\sum_{j=1}^{k_{2}}X_{i,j}\leq 0$
@@
To prove that initial $\Delta^{0}$ satisfies $\mathcal{P}$, we list three possibilities on $k_{1},k_{2}$.
@@
    if $k_{1}=k_{2}=k$, then by the block constraint $\sum_{i=1}^{k}\sum_{j=1}^{k}X_{i,j}\leq k-\gamma_{k}$, we have
    $\sum_{i=1}^{k}\sum_{j=1}^{k}\Delta^{0}_{i,j}\leq 0$
@@
    if $k_{1}>k_{2}$, then consider the first $k_{2}$ column constraints, by feasibility of $\Delta^{0}$,we have for every $1\leq i\leq k_{2}$,
@@
    if $k_{1}<k_{2}$, similar to the second case, just swapping the column and row.
@@
    Last but not least, $\mathcal{P}$ preserves $\mathcal{O}$. The only nontrivial case is when the block is specified by $(k_{1},k_{2})=(l_{1},l_{2})$. In that case, notice that by the choice of $(l_{1},l_{2})$, we have $\Delta^{0}_{i,j}=0$ when $i+j<l_{1}+l_{2}=k_{1}+k_{2}$,
    $\sum_{i=1}^{k_{1}}\sum_{j=1}^{k_{2}}\Delta^{0}_{i,j}=0$
B.8 Optimality of Lemma B.1
In this subsection, we argue that lemma B.1 is optimal in the sense that the right-hand side can not be improved in the order of dimension $p$.
Proposition B.7.
@@
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}\geq\frac{1}{2p}$
@@
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}\geq\frac{1}{\|\Delta\|_{0}}=\frac{1}{2p}$
Appendix C Proofs in Section 5
C.1 Proof of Corollary 5.1

### v2

@@ D.1 Proof sketch
    Let $g:=\sqrt{\frac{p}{\mu r}}|\lambda|$, by considering separately the cases of $r=1$ and $r\geq 2$, it is straightforward to show
    $(r-1)\cdot\big{(}\frac{1}{2r}\sqrt{\frac{p}{\mu r}}|\lambda|\big{)}^{2}\leq\frac{g^{2}}{8}.$
    Putting together, we get
    $\|V^{\top}Re_{1}\|^{2}$ $\leq(1-\frac{g}{4})^{2}+\frac{g^{2}}{8}$
    $\leq 1-\frac{g}{2}+\frac{3g^{2}}{16}\leq 1-\frac{g}{8}.$ (52)
@@
    Hence by the same technique as (41), and by (D.4), we have
@@
    $\geq 1-\sqrt{1-\frac{g}{8}}.$
    Further, by the definition of $g$ and $|\lambda|\leq 2\sqrt{\frac{\mu r}{p}}$,
    $1-\sqrt{1-\frac{g}{8}}$ $\geq\frac{g}{16}=\frac{1}{16}\sqrt{\frac{p}{\mu r}}|\lambda|\geq\frac{a^{2}|\lambda|^{2}}{32}\frac{p}{\mu r^{2}}.$
    Therefore, it holds that $r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{a^{2}|\lambda|^{2}}{32}\frac{p}{\mu r^{2}}$.
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq ca^{2}|\lambda|^{2}\frac{p}{\mu r^{2}}=c\frac{p}{\mu r^{2}}a^{2}(r_{1}-v_{1})^{2}.$
D.5 Proof of Lemma D.5
@@
We argue that there is no loss of generality in assuming $u=ae_{1}$ and $l=be_{1}$. To see that, notice that the lemma is invariant up to joint rotation. Specifically, let
$\widetilde{U}=US,\widetilde{V}=VS^{\top},\widetilde{L}=LT,\widetilde{R}=RT^{\top}.$
@@
where $S,T\in\mathcal{O}_{r}$ are arbitrary orthogonal matrices. It follows that
$\langle U^{\top}L,V^{\top}R\rangle=\langle\widetilde{U}^{\top}\widetilde{L},\widetilde{V}^{\top}\widetilde{R}\rangle,u^{\top}v=\widetilde{u}^{\top}\widetilde{v},l^{\top}r=\widetilde{l}^{\top}\widetilde{r}.$
In addition, the incoherence parameter $\mu$ stays unchanged after the procedure. Thus, we can rotate these matrices such that the first row of $U$ and $L$ align with $e_{1}$ while preserving the lemma statement, i.e., $u=ae_{1}$ and $l=be_{1}$. Furthermore, we can assume $b\geq a\geq 0$, as $\langle U^{\top}L,V^{\top}R\rangle=\langle L^{\top}U,R^{\top}V\rangle$ and we can swap the order of $U$ and $L$, $V$ and $R$.
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq c\frac{p}{\mu r^{2}}\cdot(av_{1}-br_{1})^{2}.$
@@
Now by Lemma D.3 we have
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{2\mu r}\cdot r_{1}^{2}(a-b)^{2}\geq\frac{p}{32\mu r^{2}}r_{1}^{2}(a-b)^{2}.$
By Lemma D.4 we have
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{32\mu r^{2}}a^{2}(r_{1}-v_{1})^{2}.$
@@
$r-\langle U^{\top}L,V^{\top}R\rangle$ $\geq\frac{1}{2}\cdot\frac{p}{32\mu r^{2}}\big{(}r_{1}^{2}(a-b)^{2}+a^{2}(r_{1}-v_{1})^{2}\big{)}$
$\geq\frac{p}{128\mu r^{2}}(av_{1}-br_{1})^{2}.$
@@
In fact, using Lemma D.2 with $r=1$. We get
$U^{\top}L\leq 1-\frac{(u_{1}-l_{1})^{2}}{2}$
@@
$\geq 1-\sqrt{1-\frac{(u_{1}-l_{1})^{2}}{2}}\geq\frac{(u_{1}-l_{1})^{2}}{4}.$
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{(u_{1}-l_{1})^{2}}{4}\geq\frac{p}{\mu r}\frac{v_{1}^{2}(u_{1}-l_{1})^{2}}{4}.$
@@
$r-\langle U^{\top}L,V^{\top}R\rangle\geq\frac{p}{\mu r}\frac{l_{1}^{2}(v_{1}-r_{1})^{2}}{4}.$
@@
$\geq\frac{p}{\mu r^{2}}\frac{(u_{1}v_{1}-l_{1}r_{1})^{2}}{16}.$
D.6 Proof of Lemma D.1
We postpone the technical Lemma D.6 to the next subsection and state the proof of the key separation lemma.
Proof of lemma D.1.
@@
$=\sum_{i=1}^{r}\sigma_{i}^{2}+\sum_{i=1}^{r}\lambda_{i}^{2}-2\operatorname{Tr}(R^{\top}V\Sigma U^{\top}L\Lambda).$
@@
Note that Lemma D.5 holds if we replace $r$ by $1\leq k\leq r$. In specific,
$k-\langle U_{[1:k]}^{\top}L_{[1:k]},V_{[1:k]}^{\top}R_{[1:k]}\rangle\geq c\frac{p}{\mu k^{2}}\big{[}\sum_{i=1}^{k}(u_{i}v_{i}-l_{i}r_{i})\big{]}^{2}\geq c\frac{p}{\mu r^{2}}\big{[}\sum_{i=1}^{k}(u_{i}v_{i}-l_{i}r_{i})\big{]}^{2}.$
Hence it’s not hard to draw the conclusion that $X\in\Omega$ where $\Omega$ is defined in (35) with
$\gamma_{k}=c\frac{p}{\mu r^{2}}\big{[}\sum_{i=1}^{k}(u_{i}v_{i}-l_{i}r_{i})\big{]}^{2}.$ (53)
@@
$\gamma_{k}\leq c\frac{p}{\mu r^{2}}2r\frac{\mu r}{p}=2c<\frac{1}{3}.$
We can apply Lemma D.6 and so
@@
$\geq\sum_{i=1}^{r}\frac{1}{3}(\sigma_{i}-\lambda_{i})^{2}+\sum_{i=1}^{r-1}\gamma_{i}\{(\sigma_{i}-\lambda_{i+1})^{2}+(\lambda_{i}-\sigma_{i+1})^{2}\}+\gamma_{r}(\sigma_{r}^{2}+\lambda_{r}^{2}),$ (54)
where $A$ and $X_{0}$ are defined in Lemma D.6, and the second equality follows from the direct calculation, and we let $\gamma_{0}$ for convenience.
On the other hand, let $\phi_{i}=\left\{\begin{aligned} &\sigma_{i},\quad\text{if}\;2\ |\ i\\ &\lambda_{i},\quad\text{if}\ 2\nmid i\end{aligned}\right.$, we have by Cauchy-Schwarz inequality,
$\Delta_{1,1}^{2}=(P_{1,1}-Q_{1,1})^{2}$ $=\Big{[}\sum_{i=1}^{r}\big{(}\sigma_{i}u_{i}v_{i}-\lambda_{i}l_{i}r_{i}\big{)}\Big{]}^{2}$
$\leq 2\Big{[}\sum_{i=1}^{r}\big{(}\phi_{i}u_{i}v_{i}-\phi_{i}l_{i}r_{i}\big{)}\Big{]}^{2}+2r\sum_{i=1}^{r}(\sigma_{i}-\lambda_{i})^{2}(\frac{\mu r}{p})^{2}.$ (55)
The first term in the above display can be rewritten as
@@
$\Big{[}\sum_{i=1}^{r}\big{(}\phi_{i}u_{i}v_{i}-\phi_{i}l_{i}r_{i}\big{)}\Big{]}^{2}$ $\leq r\sum_{k=1}^{r}(\phi_{k}-\phi_{k+1})^{2}\Big{[}\sum_{i=1}^{k}(u_{i}v_{i}-l_{i}r_{i})\Big{]}^{2}$
$=r\frac{\mu r^{2}}{cp}\sum_{k=1}^{r}(\phi_{k}-\phi_{k+1})^{2}\gamma_{k}$
$\leq\frac{\mu r^{3}}{cp}\sum_{i=1}^{r-1}\{(\sigma_{i}-\lambda_{i+1})^{2}+(\lambda_{i}-\sigma_{i+1})^{2}\}^{2}\gamma_{i}+\frac{\mu r^{3}}{cp}(\sigma_{r}^{2}+\lambda_{r}^{2})\gamma_{r},$ (56)
where the first inequality uses Cauchy-Schwarz inequality, the first equality uses the definition (53), and the last inequality follows from the definition of $\phi$.
Therefore, choosing $c$ sufficiently small and combining (54),(55),(56) gives
$\Delta_{1,1}^{2}$ $\leq\frac{2\mu r^{3}}{cp}\sum_{i=1}^{r-1}\{(\sigma_{i}-\lambda_{i+1})^{2}+(\lambda_{i}-\sigma_{i+1})^{2}\}^{2}\gamma_{i}+\frac{2\mu r^{3}}{cp}(\sigma_{r}^{2}+\lambda_{r}^{2})\gamma_{r}+2r\sum_{i=1}^{r}(\sigma_{i}-\lambda_{i})^{2}(\frac{\mu r}{p})^{2}$
$\leq\frac{2\mu r^{3}}{cp}\left(\sum_{i=1}^{r}\frac{1}{3}(\sigma_{i}-\lambda_{i})^{2}+\sum_{i=1}^{r-1}\gamma_{i}\{(\sigma_{i}-\lambda_{i+1})^{2}+(\lambda_{i}-\sigma_{i+1})^{2}\}+\gamma_{r}(\sigma_{r}^{2}+\lambda_{r}^{2})\right)$
$\leq\frac{2\mu r^{3}}{cp}\|\Delta\|_{F}^{2},$
where the second inequality utilizes implicitly $p\geq 3\mu c$. In fact, for $c\leq\frac{1}{2}$, if $p<3\mu c$, it holds that $\frac{2\mu r^{3}}{cp}\geq 1$, rendering (32) vacuous. Set $\tilde{c}=\frac{2}{c}$ finishes the proof.
D.7 Proof of Lemma D.6
Proof of lemma D.6.
@@
    It is straightforward to observe that $\Omega$ is a polyhedron, and the problem is LP.
@@
    Because the $\epsilon$-ball $\{X\big{|}\|X\|_{F}\leq\epsilon\}\subset\Omega$ if we choose $\epsilon$ small enough, we can deduce that $\Omega$ is non-degenerate.
@@
Finally, we prove that moving in every feasible direction from $X_{0}$ will not increase the value of the objective function, hence finishing the proof.
To obtain that, let $\Delta^{0}$ be an arbitrary feasible direction, and we want to prove $\langle A,\Delta^{0}\rangle\leq 0$ where $A$ is defined in (34). In later proof, we construct an operation ${\cal O}$ and a property ${\cal P}$ such that the initial $\Delta^{0}$ satisfies the property, the operation will not decrease the objective function, i.e., $\langle A,{\cal O}(\Delta^{0})\rangle\geq\langle A,\Delta^{0}\rangle$ while preserving the prescribed property ${\cal P}$, and the operations will end up being a zero matrix in finite steps. Finally, we have ${\cal O}^{I}(\Delta^{0})$ will end up being a zero matrix for some integer $I$, where ${\cal O}^{I}$ is the order-$I$ composition of ${\cal O}$. And hence $\langle A,\Delta^{0}\rangle\leq\langle A,{\cal O}^{I}(\Delta^{0})\rangle=0$.
The property $\mathcal{P}$ defined for a $X\in\mathbb{R}^{r\times r}$ is sum of every upper left block of $X$ is greater or equal to 0, that is, for every $1\leq k_{1},k_{2}\leq r$, it requires
$\sum_{i=1}^{k_{1}}\sum_{j=1}^{k_{2}}X_{i,j}\leq 0.$
@@
Denote the smallest sum of row index and column index for a nonzero element of $X$ as ${\cal I}(X)$. It can be seen that ${\cal I}(X)<{\cal I}({\cal O}(X))$when $X\neq 0$, which implies that the iterative operations on $\Delta^{0}$ will end up being a zero matrix in finite steps. To prove that initial $\Delta^{0}$ satisfies $\mathcal{P}$, we list three possibilities on $k_{1},k_{2}$.
@@
    if $k_{1}=k_{2}=k$, then by the block constraint $\sum_{i=1}^{k}\sum_{j=1}^{k}X_{i,j}\leq k-\gamma_{k}$ and that this constraint is tight on $X_{0}$, we have
    $\sum_{i=1}^{k}\sum_{j=1}^{k}\Delta^{0}_{i,j}\leq 0.$
@@
    if $k_{1}>k_{2}$, then consider the first $k_{2}$ column constraints, by feasibility of the direction $\Delta^{0}$ and the tightness of these constraints on $X_{0}$, we have for every $1\leq i\leq k_{2}$,
@@
    if $k_{1}<k_{2}$, the proof is similar to the second case by swapping the column and row.
@@
    Last but not least, $\mathcal{O}$ preserves $\mathcal{P}$. The only nontrivial case is when the block is specified by $(k_{1},k_{2})=(l_{1},l_{2})$. In that case, notice that by the choice of $(l_{1},l_{2})$, we have $\Delta^{0}_{i,j}=0$ when $i+j<l_{1}+l_{2}=k_{1}+k_{2}$, hence ${\cal O}(\Delta^{0})_{i,j}=0$ when $i+j<l_{1}+l_{2}=k_{1}+k_{2}$, and
    $\sum_{i=1}^{k_{1}}\sum_{j=1}^{k_{2}}{\cal O}(\Delta^{0})_{i,j}=0.$
D.8 Optimality of Lemma D.1
In this subsection, we demonstrate that Lemma D.1 is optimal in the sense that the right-hand side can not be improved in the order of dimension $p$, the incoherence parameter $\mu$. Moreover, it is expected to grow at least linearly with the rank $r$, though the precise relationship remains undetermined. To begin, we present a simplified version focusing solely on the dimension.
Proposition D.7.
@@
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}\geq\frac{1}{2p}.$
@@
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}\geq\frac{1}{\|\Delta\|_{0}}=\frac{1}{2p}.$
We can prove a more refined version below.
Proposition D.8.
For any $\mu>1,r\geq 1$ such that $\mu r<p$, there exists two rank-$r$ and $\mu$-incoherent matrices $P$ and $Q$ such that for $\Delta=P-Q$, it holds that
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}\geq\frac{\mu r}{2p}.$
Remark 19.
The lower bound in Proposition D.8 scales with $r$ instead of $r^{3}$ in Lemma D.1, the optimal scaling with respect to $r$ is left to future work.
Proof.
Let $\tilde{U}\in{\cal O}_{p,r}$ be a $(1+\mu)/2$-incoherent orthogonal matrix such that $\|\tilde{U}^{\top}e_{1}\|\in[\frac{1}{2}\sqrt{\frac{\mu r}{p}},\frac{1}{2}]$. Denote $a=\|\tilde{U}^{\top}e_{1}\|$. Let $Q_{1}\in{\cal O}_{r}$ such that $Q_{1}\tilde{U}=ae_{1}$. Denote $U=Q_{1}\tilde{U}$, then we also have $U\in{\cal O}_{p,r}$.
Denote $Ue_{1}=\begin{bmatrix}a\\ \tilde{u}\end{bmatrix}$, then $\|\tilde{u}\|=\sqrt{1-a^{2}}$. Now we start to define $V$. Pick $b$ such that $a<b<\min\{\sqrt{\frac{\mu r}{p}},\frac{1}{2}\}$. We then set $Ve_{1}=\begin{bmatrix}b\\ \tilde{v}\end{bmatrix}$, where $\tilde{v}=\sqrt{\frac{1-b^{2}}{1-a^{2}}}\tilde{u}$, and $Ve_{i}=Ue_{i}$ for $2\leq i\leq r$. It is easy to verify that $U,V\in{\cal O}_{p,r}$ and are both $\mu$-incoherent.
Let $P=UU^{\top}$ and $Q=VV^{\top}$. We have that
$\|U^{\top}V\|_{F}^{2}$ $=(V^{\top}Ue_{1})^{2}+\sum_{i=2}^{r}(V^{\top}Ue_{i})^{2}$
$=r-1+\Big{(}ab+\sqrt{(1-a^{2})(1-b^{2})}\Big{)}^{2}$
We have that
$\|\Delta\|_{F}^{2}$ $=r-\langle U^{\top}V,U^{\top}V\rangle$
$=r-\|U^{\top}V\|_{F}^{2}$
$=1-\Big{(}ab+\sqrt{(1-a^{2})(1-b^{2})}\Big{)}^{2}$
$=\Big{(}a\sqrt{1-b^{2}}-b\sqrt{1-a^{2}}\Big{)}^{2}$
$=\frac{(a-b)^{2}(a+b)^{2}}{\Big{(}a\sqrt{1-b^{2}}+b\sqrt{1-a^{2}}\Big{)}^{2}}$
$\leq 2(a-b)^{2},$
where we used $a,b\leq\frac{1}{2}$ in the inequality.
On the other hand, we have
$\|\Delta\|_{\max}\geq|e_{1}^{\top}\Delta e_{1}|=|a^{2}-b^{2}|.$
Hence
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}$ $\geq\frac{(a^{2}-b^{2})^{2}}{2(a-b)^{2}}$
$=\frac{(a+b)^{2}}{2}\geq\frac{\mu r}{2p},$
where we use that $b>a>\frac{1}{2}\sqrt{\frac{\mu r}{p}}$.
Appendix E Proofs in Section 4
E.1 Proof of Corollary 4.1